Solve the following problems.show the equation used for the solution.
length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 34 in., find the dimensions of the rectangle.

Answer by ptaylor(2198)

Let w=width of the rectangle
Then length(l)=2w+2
Perimeter=2l+2w so our equation to solve is:
34=2(2w+2)+2w get rid of parens
34=4w+4+2w subtract 4 from both sides
34-4=4w+4-4+2w collect like terms
30=6w divide both sides by 6
w=5in-------------------------------width of rectangle
2w+2=2*5+2=12in--------------------length of rectangle
CK
34=2*5+2*12
34=10+24
34=34
also
12=2*5+2
12=12

Hope this helps---ptaylor